OpenStudy (anonymous):

i need someone that is good in math will fan and give medal State whether either extreme in the data set is an outlier. A. The upper extreme is an outlier. B. The lower extreme is an outlier. C. Both extremes are outliers. D. Neither extreme is an outlier.

OpenStudy (anonymous):

it looks like the answer would be c.

OpenStudy (anonymous):

thx could you help me with 4 more

OpenStudy (anonymous):

sure ok

OpenStudy (anonymous):

are you a student

OpenStudy (anonymous):

yes i am

OpenStudy (anonymous):

State whether each extreme in the data set is an outlier. A. The lower extreme is an outlier, but the upper extreme is not. B. The upper extreme is an outlier, but the lower extreme is not. C. Both extremes are outliers. D. Neither extreme is an outlier.

OpenStudy (anonymous):

what grade i am in 6th

OpenStudy (anonymous):

i am an 11th grader

OpenStudy (anonymous):

i think the answer is

OpenStudy (anonymous):

Bee

OpenStudy (anonymous):

what about you

OpenStudy (anonymous):

You have to find interquartile range, which is Q3 - Q1 Q3 = 41 Q1 = 36 IQR = Q3 - Q1 = 41 - 36 = 5 The outliers will be any values below 36 - 1.5 * 5 = 36 - 7.5 = 28.5 or above 41 + 1.5 * 5 = 41 + 7.5 = 48.5 Therefore, your outliers are : 25 and 45 (C)

OpenStudy (anonymous):

i have done this queston before in school

OpenStudy (anonymous):

For the data in the table, the mean score is 19 and the median score is 18. How are the mean and median affected when the outlier 1 is added to the data? Table titled Points Scored with data in the first and only row being 16, 17, 18, 23, 20, 18, 15, and 25. A. The mean and median are not affected. B. The mean decreases by 2, but the median stays the same. C. The median decreases by 2, but the mean stays the same. D. The mean and median both decrease by 2.

OpenStudy (anonymous):

is it confusing

OpenStudy (anonymous):

this one is b. Mean without outlier: 19 Median without outlier: 18 Mean with outlier: 17 Median with outlier: 18

OpenStudy (anonymous):

thx 32 more

OpenStudy (anonymous):

i mean 2 more

OpenStudy (anonymous):

lol woah i was about to say 32 is alot

OpenStudy (anonymous):

Are the data shown in the line plot skewed left, skewed right, or not skewed? A. skewed left B. skewed right C. not skewed

OpenStudy (anonymous):

so this one took me awhile but it is A. skewed left

OpenStudy (anonymous):

The data shown are hourly wages of some employees of a small company. $5, $7, $8, $10, $24, $40 Which values, if any, are outliers? A. $5 only B. $40 only C. both $5 and $40 D. none

OpenStudy (anonymous):

this is a test and i have 5 questions this is the last one

OpenStudy (anonymous):

@Siri15 you left why

OpenStudy (anonymous):

sorry i had to do the math but this on is D.none

OpenStudy (anonymous):

let;s see what i got

OpenStudy (anonymous):

man i got a 60%

OpenStudy (anonymous):

i could retake it but what if i get alower score

OpenStudy (anonymous):

thats weird all my calculations were correct my professor even checked them

OpenStudy (anonymous):

well my computer is maybe crazy

OpenStudy (anonymous):

though i have to get an 80 or above to pass

OpenStudy (anonymous):

thx anyway you helped

OpenStudy (anonymous):

well in my case i work on an cumputer 8 hours a day

OpenStudy (anonymous):

and your welcome

OpenStudy (anonymous):

me too i am homeschooled so i stay all day seated

OpenStudy (anonymous):

well have a nice day

OpenStudy (anonymous):

i go to school but not a ordinary school i go to a school were i graduate early and im suppose to graduate in December but i came in as a 9th grader but since im going to graduate in December im known as a 11th grader

OpenStudy (anonymous):

WOW

OpenStudy (anonymous):

have a nice day

OpenStudy (anonymous):

lol you too